According to the reaction C2H6O(g) + O2(g) = CO2(g) + H2O(g). 31.7 grams of C2H6O reacts eith 23.5 grams of oxygen to produce ____ grams of water vapor. What is yhr percent yield of the rreaction if 30.0 grams of carbon dioxide were actually made?


1
Expert's answer
2020-11-04T14:22:42-0500

Balanced equation : C2H6O +3O2 = 2CO2 +3H2O

The limiting reagent in this reaction is O2

3O2 =3H2O

[3×16×2g] O2 = (3×18)g H20

1g O2= (3×18)/(3×32)g H20

23.5g O2= (3×18)/(3×32)×23.5g H20

23.5g O2 will produce 13.22g of H2O


3O2= 2CO2

3×32g O2= 2× 44g CO2

1g O2= (2× 44)/(3×32)g CO2

23.5g O2= (2× 44)/(3×32)× 23.5g CO2

23.5g O2 will produce 21.54g CO2

% yield =[actual yield / theoretical yield] ×100%

%yield = [30/21.54] ×100% = 139.28%


Usually the actual yield < theoretical yield.

So, there seem to be an error in the data provided